# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v4

@@ III PROBLEM FORMULATION
We consider a spherical robot with radius $r\geq 0$ operating in a convex, three-dimensional Euclidean space $\mathcal{W}\subseteq\mathbb{R}^{3}.$ The workspace $\mathcal{W}$ is cluttered with finite number of compact, convex obstacles $\mathcal{O}_{i}\subset\mathcal{W},i\in\{1,\ldots,b\},b\in\mathbb{N}.$ We define obstacle $\mathcal{O}_{0}:=(\mathcal{W}^{\circ})^{c}$ as the complement of the interior of the workspace. Collectively, $\mathcal{O}_{\mathcal{W}}=\bigcup_{i\in\mathbb{I}}\mathcal{O}_{i}$, where $\mathbb{I}=\{0,\ldots,b\}.$ We make the following workspace feasibility assumption:
@@
$d(\mathcal{O}_{i},\mathcal{O}_{j}):=\underset{\mathbf{p}\in\mathcal{O}_{i},\mathbf{q}\in\mathcal{O}_{j}}{\text{min}}\|{\mathbf{p}-\mathbf{q}}\|\geq 2(r+\delta).$ (8)
We then pick an arbitrary value $r_{s}\in(0,\delta)$ as the minimum distance that the robot should maintain with respect to any obstacle. The obstacle-free workspace is then defined as $\mathcal{W}_{0}:=\mathcal{W}\setminus\mathcal{O}_{\mathcal{W}}.$ Given $y\geq 0$, a $y-$eroded obstacle-free workspace, $\mathcal{W}_{y}$ is defined as
$\mathcal{W}_{y}:=\mathcal{W}\setminus\bigcup_{i\in\mathbb{I}}\mathcal{D}_{y}^{\circ}(\mathcal{O}_{i})\subseteq\mathcal{W}_{0}.$ (9)
Hence, $\mathcal{W}_{r_{a}}$ with $r_{a}=r+r_{s}$ is the obstacle-free workspace with respect to the center of the robot i.e., $\mathbf{x}\in\mathcal{W}_{r_{a}}\iff\mathcal{B}_{r_{a}}(\mathbf{x})\subset\mathcal{W}_{0}.$
@@
where $\mathbf{u}\in\mathbb{R}^{3}$ is the control input. The task is to design a feedback control law $\mathbf{u}$ such that:
@@
    Safety: the obstacle-free workspace $\mathcal{W}_{r_{a}}$ is forward invariant with respect to the center of the robot,
@@
    Global Asymptotic Stability: Any given target location $\mathbf{x}^{\mathbf{d}}\in(\mathcal{W}_{r_{a}})^{\circ}$ is a globally asymptotically stable equilibrium for the closed-loop system. Without loss of generality, we will consider $\mathbf{x}^{\mathbf{d}}=\mathbf{0}$.
@@ IV HYBRID CONTROL FOR OBSTACLE AVOIDANCE
In the proposed scheme, the robot operates in two modes based on the value of a mode indicator variable $m\in\mathbb{M}:=\{0,1\}$. The move-to-target mode $(m=0)$ is adopted when the robot is away from the obstacles and the obstacle-avoidance mode $(m=1)$ is adopted when the robot is close to an obstacle obstructing its motion towards the target location. In the move-to-target mode, the robot moves straight towards the target location. In the obstacle-avoidance mode, the robot navigates around the obstacle while staying within its $\gamma-$neighborhood, where $\gamma\in(0,\delta-r_{s})$. When the robot operates in the obstacle-avoidance mode, the robot’s center has a unique closest point on the nearest obstacle. Furthermore, to prevent the robot from getting trapped in a loop around an obstacle, the proposed obstacle-avoidance strategy confines the robot’s center to a hyperplane that passes through the target location. It then guides the robot along the boundary to a position where the nearest obstacle no longer intersects with the line segment connecting the robot’s center and the target location, as shown in Fig. 1a.
@@ IV-A Hybrid control design
$\mathbf{u}(\xi)=-\kappa_{s}(1-m)\mathbf{x}+\kappa_{r}m\mathbf{v}(\mathbf{x},\mathbf{a}),$ (11a)
$\underbrace{\begin{matrix}\mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}\;=\mathbf{0},\\ \;=0,\\ \;=0\\ \;=1,\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{F}}\quad\quad\underbrace{\begin{bmatrix}\mathbf{h^{+}}\\ \mathbf{a^{+}}\\ m^{+}\\ {s^{+}}\end{bmatrix}\begin{matrix}\;\in\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s),\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{J}}$ (11b)
@@
where $\kappa_{s}>0$, $\kappa_{r}>0$, and $\xi=(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{K}:=\mathcal{W}_{r_{a}}\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{2}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$. The variable $\mathbf{h}$ denotes the hit point, which is the location where the robot switches from the move-to-target mode to the obstacle-avoidance mode. The vector $\mathbf{a}\in\mathbb{S}^{2}$ is instrumental for the construction of the avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{a})$, used in (11a). The scalar variable $s\in\mathbb{R}_{\geq 0}$ allows the robot to switch from the obstacle-avoidance mode to the move-to-target mode when initially in the obstacle-avoidance mode. Details of this switching process are provided later in Section IV-B. The sets $\mathcal{F}$ and $\mathcal{J}$ are the flow and jump sets related to different modes of operation, respectively, whose constructions are provided in Section IV-B. The update law $\mathbf{L}$, which allows the robot to update the values of the variables $\mathbf{h}$, $\mathbf{a}$, $m$ and $s$ based on the current location of the robot with respect to the nearest obstacle and the target location, will be designed later in Section IV-C. Next, we provide the design of the vector $\mathbf{v}(\mathbf{x},\mathbf{a})\in\mathbb{R}^{3}$.
The vector $\mathbf{v}(\mathbf{x},\mathbf{a}),$ used in (11a), is defined as
$\mathbf{v}(\mathbf{x},\mathbf{a})=\big{[}\eta(\mathbf{x})\mathbf{I}_{3}+(1-|\eta(\mathbf{x})|)\mathbf{R}(\mathbf{a})\big{]}{{\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}}}$ (12)
@@
where $\mathbf{I}_{3}\in\mathbb{R}^{3\times 3}$ is the identity matrix and $\mathbf{x}_{\pi}:=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}).$ The location $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is the closest point on the set $\mathcal{O}_{\mathcal{W}}$ from the robot’s center $\mathbf{x}$, as defined in Section II-B. Notice that, since the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$ are convex and the parameter $\gamma\in(0,\delta-r_{s})$, according to Assumption 8, the robot will have a unique closest point to the obstacles whenever its center is in the $(r_{a}+\gamma)-$neighborhood of these obstacles. On the other hand, there may be some locations in the $(r_{a}+\gamma)-$neighborhood of the obstacle $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$ for which the uniqueness of the closest point from robot’s center to the obstacle $\mathcal{O}_{0}$ cannot be guaranteed. However, as discussed later in Remark 1, the design of the flow sets and jump sets guarantees that the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{a})$ is never activated in the region $\mathcal{N}_{r_{a}+\gamma}(\mathcal{O}_{0}).$
The rotation matrix $\mathbf{R}(\mathbf{a}):=\mathcal{R}(\frac{\pi}{2},\mathbf{a})\in SO(3)$, with $\mathcal{R}(\theta,\mathbf{a})$ being the rotation by an angle $\theta$ about the unit vector $\mathbf{a}$, defined as follows:
$\mathcal{R}(\theta,\mathbf{a}):=\mathbf{I}_{3}+\sin(\theta)\mathbf{a}^{\times}+(1-\cos(\theta))(\mathbf{a}^{\times})^{2},$
which, for $\theta=\frac{\pi}{2}$, leads to
$\mathbf{R}(\mathbf{a})=\mathbf{a}\mathbf{a}^{\top}+\mathbf{a}^{\times},$
where we used the fact that $(\mathbf{a}^{\times})^{2}=\mathbf{a}\mathbf{a}^{\top}-\mathbf{I}_{3}$, with $\mathbf{a}^{\times}$ being the skew-symmetric matrix associated to the unit vector $\mathbf{a}=[a_{1}~{}a_{2}~{}a_{3}]^{\top}$, defined according to (1).
The matrix $\mathbf{P}(\mathbf{a})\in\mathbb{R}^{3\times 3}$, which is used in (12), is given by
$\mathbf{P}(\mathbf{a}):=\mathbf{I}_{3}-{\mathbf{a}\mathbf{a}^{\intercal}},$ (13)
@@
where $\mathbf{a}\in\mathbb{S}^{2}$. For any vector $\mathbf{x}\in\mathbb{R}^{3},$ the vector $\mathbf{P}(\mathbf{a})\mathbf{x}$ corresponds to the projection of $\mathbf{x}$ onto the hyperplane orthogonal to $\mathbf{a}$. As discussed later in Section IV-C, the coordinates of the unit vector $\mathbf{a}$ are updated when the robot switches from the move-to-target mode to the obstacle-avoidance mode according to the update law $\mathbf{L}(\xi)$, whose design is provided later in Section IV-C.
@@
$\eta(\mathbf{x})=\begin{cases}-1,&d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}\geq\gamma_{s},\\ 1-\frac{d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}-\gamma_{a}}{0.5(\gamma_{s}-\gamma_{a})},&\gamma_{a}\leq d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}\leq\gamma_{s},\\ 1,&d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}\leq\gamma_{a},\end{cases}$ (14)
where $0<\gamma_{a}<\gamma_{s}<\gamma.$ The scalar function $\eta$ is designed to ensure that the center of the robot remains inside the $\gamma-$neighborhood of the $r_{a}-$dilated obstacle-occupied workspace $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$ when it operates in the obstacle-avoidance mode in the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$. This feature allows for the design of the jump set of the obstacle-avoidance mode, as discussed later in Section IV-B2, and ensures convergence to the target location, as stated later in Theorem 1. Next, we provide the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$ used in (11).
Figure 1: (a) Robot trajectory moving along the hyperplane $\mathcal{P}(\mathbf{h},\mathbf{a})$ while avoiding obstacle $\mathcal{O}_{i}$. (b) The geometrical representation of the landing region and the exit region.

### v5

@@ III Problem Formulation
Let $\mathcal{W}$ be a closed subset of the $n-$dimensional Euclidean space that bounds the workspace. The workspace $\mathcal{W}$ consists of finite number of compact, convex obstacles $\mathcal{O}_{i}$, $i\in\{1,\ldots,b\},b\in\mathbb{N}$. We define obstacle $\mathcal{O}_{0}:=(\mathcal{W}^{\circ})^{c}$ as the complement of the interior of the workspace. Collectively, the obstacle-occupied workspace is denoted by $\mathcal{O}_{\mathcal{W}}=\bigcup_{i\in\mathbb{I}}\mathcal{O}_{i}$, where $\mathbb{I}=\{0,\ldots,b\}.$
The robot is represented by a $n-$dimensional sphere with radius $r\geq 0$ and center $\mathbf{x}$. It is equipped with a range sensor that can identify the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ which contains the locations on the boundaries of nearby obstacles, provided there is a clear line of sight to the center of the robot, up to a certain sensing range $R_{s}$, as shown in Fig. 1. The set ð$\eth_{\mathcal{W}}$ is defined as follows:
ð$\eth_{\mathcal{W}}^{\mathbf{x}}:=\{\mathbf{p}\in\partial\mathcal{O}_{\mathcal{W}}|\|\mathbf{p}-\mathbf{x}\|\leq R_{s},\mathcal{L}_{s}(\mathbf{p},\mathbf{x})\cap\mathcal{O}_{\mathcal{W}}=\mathbf{p}\}.$ (8)
Figure 1: Illustration of the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ as defined in (8).
To maintain the local convexity of the obstacle boundaries, we impose the following workspace feasibility assumption.
@@
$d(\mathcal{O}_{i},\mathcal{O}_{j}):=\underset{\mathbf{p}\in\mathcal{O}_{i},\mathbf{q}\in\mathcal{O}_{j}}{\text{min}}\|{\mathbf{p}-\mathbf{q}}\|\geq 2(r+\delta),$ (9)
where $\delta>0$.
We then pick an arbitrarily small value $r_{s}\in(0,\delta)$ as the minimum distance that the robot should maintain to any obstacle. Therefore, the obstacle-free workspace with respect to the center of the robot is defined as follows:
$\mathcal{W}_{r_{a}}:=\mathcal{W}\setminus(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))^{\circ},$ (10)
where $r_{a}=r+r_{s}$. According to Assumption 9, the set $\mathcal{W}_{r_{a}}$ is pathwise connected and $\mathbf{x}\in\mathcal{W}_{r_{a}}\iff\mathcal{B}_{r_{a}}(\mathbf{x})\subset\mathcal{W}\setminus\mathcal{O}_{\mathcal{W}}^{\circ}$.
@@
where $\mathbf{u}\in\mathbb{R}^{n}$ is the control input. Given a workspace that satisfies Assumption 9, and given that the robot can identify the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ (8), the task is to design a feedback control law $\mathbf{u}$ to guarantee the following properties:
@@
    Safety: the obstacle-free workspace $\mathcal{W}_{r_{a}}$ with respect to the center of the robot is forward invariant,
@@
    Global asymptotic stability: any target location $\mathbf{x^{d}}\in(\mathcal{W}_{r_{a}})^{\circ}$ is a globally asymptotically stable equilibrium for the closed-loop system. Without loss of generality, we will consider $\mathbf{x^{d}}=\mathbf{0}$.
Figure 2: Robot trajectory moving along the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ while avoiding an obstacle.
@@ IV Hybrid Control For Obstacle Avoidance
The proposed hybrid feedback controller operates in two different modes based on the mode indicator $m\in\mathbb{M}:=\{0,1\}$. In the move-to-target mode $(m=0)$, the control input vector directs the robot along a straight line toward the target location. When the robot is in the neighbourhood of an obstacle that obstructs its direct path (i.e., the distance from the robot’s centre to the closest point on the obstacle is less than or equal to $\gamma\in(0,\delta-r_{s})$), the control law switches to the obstacle-avoidance mode $(m=1)$. In contrast to the two-dimensional case, where the robot can only move clockwise or counterclockwise around an obstacle, the three-dimensional environment offers an infinite number of safe paths around the obstacles. Thus, designing an obstacle-avoidance strategy that maintains the continuity of the control input vector and prevents the robot from retracing its path is crucial. The proposed strategy achieves this by confining the robot’s motion to a plane that passes through both the target location and the interior of the obstacle being avoided. The robot is guided along the obstacle’s boundary until it reaches a point where the obstacle no longer blocks the direct line-of-sight to the target location, as shown in Fig. 2.
@@ IV-A Hybrid control design
$\mathbf{u}(\xi)=-\kappa_{s}(1-m)\mathbf{x}+\kappa_{r}m\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a}),$ (12a)
$\underbrace{\begin{matrix}\mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}\;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{F}}\quad\quad\underbrace{\begin{bmatrix}\mathbf{h^{+}}\\ \mathbf{a^{+}}\\ m^{+}\\ {s^{+}}\end{bmatrix}\begin{matrix}\;\in\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s),\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{J}}$ (12b)
@@
where $\kappa_{s}>0$, $\kappa_{r}>0$, and the composite state vector $\xi:=(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{K}:=\mathcal{W}_{r_{a}}\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{n-1}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$ The variable $\mathbf{h}$ denotes the hit point, which is the location where the robot switches from the move-to-target mode to the obstacle-avoidance mode. The vector $\mathbf{a}\in\mathbb{S}^{n-1}$ is instrumental for the construction of the avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$, used in (12a). The scalar variable $s\in\mathbb{R}_{\geq 0}$ allows the robot to switch once from the obstacle-avoidance mode to the move-to-target mode only when it is initialized in the obstacle-avoidance mode. Details of this switching process are provided later in Section IV-B. The sets $\mathcal{F}$ and $\mathcal{J}$ are the flow and jump sets related to different modes of operation, respectively, whose constructions are provided in Section IV-B. The update law $\mathbf{L}$, which allows the robot to update the values of the variables $\mathbf{h}$, $\mathbf{a}$, $m$ and $s$ based on the current location of the robot with respect to the obstacle being avoided and the target location, will be designed later in Section IV-C. Next, we provide the design of the vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})\in\mathbb{R}^{n}$. The vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$, used in (12a), is defined as
$\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})=\left[\eta(\mathbf{x})+(1-|\eta(\mathbf{x})|)\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\right]\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},$ (13)
@@
where $\hat{\mathbf{h}}$ denotes unit vector in the direction of $\mathbf{h}$, and $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$. The vector $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ represents the point on the obstacle-occupied workspace closest to $\mathbf{x}$, as defined in Section II-B. Notice that, since the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$ are convex and the parameter $\gamma\in(0,\delta-r_{s})$, according to Assumption 9, the robot will have a unique closest point to the obstacles whenever its center is in the $(r_{a}+\gamma)-$neighbourhood of these obstacles. On the other hand, since obstacle $\mathcal{O}_{0}$, where $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$, is non-convex, there may be some locations in the $(r_{a}+\gamma)-$neighbourhood of the obstacle $\mathcal{O}_{0}$ for which the uniqueness of the closest point from the robot’s center to the obstacle $\mathcal{O}_{0}$ cannot be guaranteed. However, as discussed later in Remark 2, the design of the flow sets and the jump sets guarantees that the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ is never activated in the region $\mathcal{N}_{r_{a}+\gamma}(\mathcal{O}_{0}).$
@@
Note that the coordinates of the hit point $\mathbf{h}$ and the unit vector $\mathbf{a}$ are updated when the robot switches from the move-to-target mode to the obstacle-avoidance mode using the update law $\mathbf{L}(\xi)$, whose design is provided later in Section IV-C. It is ensured that $\mathbf{a}^{\top}\mathbf{h}=0$, which allows us to define the operators $\mathbf{P}$ and $\mathbf{R}$, used in (13), as discussed next.
Since $\mathbf{h}$ and $\mathbf{a}$ are ensured to be orthogonal to each other, the parallel projection operator $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})$, as described in Section II-D, is defined as
$\mathbf{P}(\hat{\mathbf{h}},\mathbf{a}):=\hat{\mathbf{h}}\hat{\mathbf{h}}^{\top}+\mathbf{a}\mathbf{a}^{\top}.$ (14)
According to (14), the vector $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$, used in (13), is the projection of the vector $\mathbf{x}_{\pi}$ onto the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$. Next, we discuss the design of the operator $\mathbf{R}$.
The rotation matrix $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a}):=R(\pi/2,\hat{\mathbf{h}},\mathbf{a})\in SO(n)$, where $R(\theta,\hat{\mathbf{h}},\mathbf{a})$ for $\theta\in[0,2\pi)$ is defined in Section II-E. For $\theta=\frac{\pi}{2}$, one gets
$\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})=\mathbf{I}_{n}+(\mathbf{a}\hat{\mathbf{h}}^{\top}-\hat{\mathbf{h}}\mathbf{a}^{\top})-(\hat{\mathbf{h}}\hat{\mathbf{h}}^{\top}+\mathbf{a}\mathbf{a}^{\top}).$ (15)
Note that, according to (14), one has $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Therefore, as per (15), the operator $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})$, used in (13), rotates the vector $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ by $\pi/2$ radians in the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$. This rotation is performed from the vector $\mathbf{h}$ to the vector $\mathbf{a}$.
@@
$\eta(\mathbf{x})=\begin{cases}-1,&d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}\geq\gamma_{s},\\ 1-\frac{d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}-\gamma_{a}}{0.5(\gamma_{s}-\gamma_{a})},&\gamma_{a}<d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}<\gamma_{s},\\ 1,&d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}\leq\gamma_{a},\end{cases}$ (16)
where $0<\gamma_{a}<\gamma_{s}<\gamma$. The scalar function $\eta$ is designed to ensure that the center of the robot remains inside the $\gamma-$neighborhood of the $r_{a}-$dilated obstacle-occupied workspace $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$ when it operates in the obstacle-avoidance mode in the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$. This feature allows for the design of the jump set of the obstacle-avoidance mode, as discussed later in Section IV-B2, and ensures convergence to the target location, as stated later in Theorem 1.
Remark 1.
Consider a plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ passing through the interior of obstacle $\mathcal{O}_{i}$ and the origin, where $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ and $\mathbf{a}\in\mathbb{S}^{n-1}$. Since $\mathcal{O}_{i}$ is a general convex obstacle, one cannot guarantee that $\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $\mathbf{x}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. Therefore, the operator $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})$ is used in (13) to ensure that $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})\in\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $\mathbf{x}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. The operator $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})$ aids in steering the robot along the boundary of obstacle $\mathcal{O}_{i}$ when it operates in the obstacle-avoidance mode. However, because $\mathcal{O}_{i}$ is a general convex obstacle, the vector $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ is not necessarily tangential to the set $\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})$ at $\mathbf{x}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, where $\beta=d(\mathbf{x},\mathcal{O}_{i})$. Therefore, it may drive the robot either closer to obstacle $\mathcal{O}_{i}$ or away from it. Thus, the scalar function $\eta(\mathbf{x})$ in included to ensure that $\mathbf{x}$ remains in the set $\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ as long as the robot operates in obstacle-avoidance mode.
Next, we provide the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$ used in (12).